OpenStudy (megannicole51):

integral of ae^((-x^2)/(5))dx=1 solve for a

OpenStudy (anonymous):

\[a \int\limits_{}^{} e^{\frac{ -x ^{2}}{ 5 }}dx = 1\] Is this right?

OpenStudy (anonymous):

Wouldn't you need to use Laplace transformations for this? (I forget how to solve it however it just seems like something I've seen before)

OpenStudy (anonymous):

What class is this? Just wondering because I might be totally off on my answer.

OpenStudy (anonymous):

@megannicole51, we need some limits for the integral.

OpenStudy (megannicole51):

there arent any

OpenStudy (megannicole51):

calc 2 at a university level

OpenStudy (megannicole51):

it says enter 3 decimal points so its not a whole number

OpenStudy (anonymous):

with no limits of integration i can't imagine how you would do this problem the integral could be almost anything also i am fairly sure that you are not going to find a nice closed form for \[\int e^{-\frac{x^2}{5}}\]

OpenStudy (loser66):

@SithsAndGiggles Can I take derivative both sides which lead to a =0? what's wrong if I do so?